[Xrayops] SR Y emittance
APS Experiment Hall X-ray Operations Problem List
xrayops at mailman.aps.anl.gov
Mon Apr 16 05:12:04 CDT 2007
On 04/16/07 at 03:00 Operaitons logged:
Storage Ring PV Monitor S:VID1:Yemittace is occasionally quickly going in and
out of alarm. It's High Limit is 0.045, is will quickly hit ~0.0457, then
quickly come back down to ~0.0394. These changes correspond with the changing
setpoints with the CPU. The PV Monitor is going in and out of alarm, but the
alarm handlers are not. We are monitoring it. Printout in the MCR.
On 04/16/07 at 04:20 Operations logged:
On 04/15/07 at 22:00 SR Y emittance was oscillating between ~0.036 to ~0.039, it
is now oscillating between ~0.039 to ~0.045. Also, the coupling is oscillating
from ~1.5% to ~1.8%. These changes correspond with the changing setpoints with
the CPU. We are continuing to monitor it.
On 04/16/07 at 04:36 Operations logged:
We have noticed the beam spot flickering, not unlike a candle in the wind. We
On 04/16/07 at 04:44 Operations logged:
We have discussed the Y emittance and beam spot flickering with Karen Schroeder.
She is investigating it from home. We are continuing to investigate it.
On 04/16/07 at 04:59 Operations logged:
Storage Ring PV Monitor S:VID1:Yemittace is now almost consistently staying in
alarm. SR Coupling is oscillating from ~1.5% and ~2.3%. The beam is still
flickering. We are continuing to investigate. The Beamline Notification PA
Announcement has been made.
On 04/16/07 at 05:10 Operations logged:
Per Karen Schroeder, we are contacting Om Singh concerning the Storage Ring. We
are continuing to investigate.
ASD Operations Group
Main Control Room
More information about the XRAYOPS